OpenStudy (anonymous):

what is the 3 solutions to this equation y=-4x-5 when x= 0,-1 and 2

OpenStudy (anonymous):

y = -4(0)-5 y = 0 - 5 y = -5 y = -4(-1) -5 y = 4 - 5 y = -1 y = -4(2) -5 y= -8 -5 y = -13

OpenStudy (anonymous):

Just substitute the values of x to the equation. :)